OpenStudy (anonymous):

Compare the graphs of the inverse variations. y=3/x and y=5/x

OpenStudy (anonymous):

y1=3/x y2=5/x y2=5/3*y1

OpenStudy (anonymous):

graphically, y2 is y1 stretched vertically by 5/3.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

:)